# Sparse Bayesian multidimensional scaling(s)

### v1

@@ 3.2 Analysis of global influenza
Holbrook et al., (2021) apply the BMDS framework to the phylogeographic analysis of the spread of influenza subtypes through transportation networks. They analyze 1,370, 1,389, 1,393 and 1,240 samples of type H1N1, H3N1, Victoria (VIC) and Yamagata (YAM), spanning 12.9, 14.2, 15.4 and 17.75 years, respectively. To scale BMDS to data of this size, they implement core model likelihood and log-likelihood gradient calculations on large graphics processing units and multi-core central processing units. Unfortunately, such an approach requires time-intensive coding and access to expensive computational hardware. We employ a similar Bayesian hierarchical model, applying the same highly structured stochastic process priors but use sBMDS to transform to a latent network space. We are interested in whether under sBMDS we can accurately and efficiently infer the subtype-specific rates of dispersal across the latent airspace for the four influenza strains.
Our data consists of pairwise “effective distances” (Brockmann and Helbing,, 2013) between countries, which inversely measures the probability of traveling between airports. More trafficked airports have a shorter “distance” and thus a higher chance of disease transmission. For strain $v$, the prior on the viral latent locations follows a multivariate Brownian diffusion process along the tree
$\mathbf{X}_{v}\sim MN(\bm{\mu}_{v},\mathbf{V}_{{\cal G}v},\mathbf{\Sigma}_{v}),$ (16)
in which $\bm{\mu}_{v}$ is the $N\times D$ mean matrix, $\mathbf{V}_{{\cal G}v}$ is the $N\times N$ row covariance matrix calculated from a fixed tree ${\cal G}v$, and $\mathbf{\Sigma}_{v}$ is the $D\times D$ column covariance matrix, independently for $v=1,...,4$. In addition, we assume a priori
@@
$d_{0}$ is the degree of freedom set as the dimension of the latent space and $\mathbf{T}_{0}$ is the rate matrix fixed as $\mathbf{I}_{D}$ in our model. We implement the adaptive HMC algorithm to recover the viral latent locations along with adaptive MH updates on the BMDS precision parameter, $1/\sigma^{2}_{v}$, and Gibbs updates on $\mathbf{\Sigma}^{-1}_{v}$. We let the latent dimension be six as Holbrook et al., (2021) recommended from 5-fold cross-validation. We find 20 leapfrog steps to be adequate as we vary the number of bands/landmarks to 50, 100 and 200. The trace of $\mathbf{\Sigma}_{v}$ provides the instantaneous rate of diffusion and is of chief scientific interest.
@@ 3.2.1 Accuracy
Figure 7 plots the posterior distributions of the strain-specific diffusion rates inferred from the full (left) and banded sparse (right) model. For each subtype and model, we run 120,000 iterations, burning the first 20,000 and saving every 100th iteration. We successfully capture the relative distributions for the B-sBMDS using 50 bands, but note that the posterior modes are slightly off. When we increase the number of bands to 200 (Figure B.3), the distributions appear identical. Using the textmineR package (Jones,, 2021) in R, we compute the Hellinger distance between the strain-specific posterior distributions of the squared effective distance per year from the full and sparse methods (Table 3). As expected, the Hellinger distance decreases with more bands.
@@ 3.2.2 Computational performance
Hellinger distance Average efficiency speedup (min, max)
@@
Hellinger distance Average efficiency speedup (min, max)
@@ 4 Discussion
We present two methods for subsetting the observed dissimilarity data: banded sparse BMDS (B-sBMDS) and landmark sparse BMDS (L-sBMDS). We show that both sparse methods obtain accurate results at low band/landmark counts even with noisy data. Moreover, combining HMC with sBMDS proves effective for inferring thousands of latent locations. Lastly, we successfully apply the sBMDS variants to four influenza subtypes using relatively low band counts and obtain diffusion rates similar to those under the much slower, full BMDS model.
Possible extensions to our work include the use of different noise distributions on the observed distances. For example, Bakker and Poole, (2013) employ Bayesian metric MDS, assuming the observed dissimilarities come from log-normal distributions. As these distributions still have $\mathcal{O}(N^{2})$ time complexity, the sBMDS could be valuable in improving the computational performance for a wider range of dissimilarity data.
Additionally, many potential theoretical developments remain. We explain in Section A.4 how one could extend Theorem 1’s proof of posterior consistency to higher dimensions. The biggest limitations are extending Lemma 3 and obtaining estimates with good dependence on dimension $D$. One could also explore treating the coupling matrix $J_{n,N}$ as a random variable that depends on the observed data (and perhaps changes over the run-time of an algorithm). An appealing feature of Raftery et al., (2012) is that they claim reasonable uncertainty quantification along a truly linear run-time. It seems difficult to formalize such a result with posterior consistency for our sBMDS models as the number of bands (landmarks) grows with the number of objects. We are left with many tantalizing questions: “by including a data-informed approach to model sparsity, can we achieve a linear run-time and still demonstrate posterior consistency?”, “how should we be measuring consistency?”, and “do the datasets Raftery et al., (2012) study have any special features that change the rate of convergence for a sBMDS-like model?”
@@ A.1 Consistent Estimates of Absolute Values
We note that $|x_{n}|$ (but not $x_{n}$ itself) is effectively identifiable given the data $\{\delta_{ny}\}_{y\in G_{n,N}^{(k)}}$, and we have the posterior concentration bound:
@@
$\mathbb{P}[p(\{u\,:\,\min(|u-x_{n}|,|u+x_{n}|)\leq c_{1}\epsilon_{N}\}|\{\delta_{ny}\}_{y\in G_{n,N}^{(k)}})\geq 1-e^{-c_{2}\ell_{N}\epsilon_{N}^{2}}]\geq 1-e^{-c_{3}\ell_{N}\epsilon_{N}^{2}}.$ (19)
@@
Given $x_{n}$, the data $\{\delta_{ny}\}_{y\in G_{n,N}^{(k)}}$ are i.i.d. with distributions being a finite mixture of truncated Gaussians. Denote the density of this distribution by $q_{x_{n}}$, and let =$\mathcal{F}=\{q_{u}\}_{u\in I}$ be the associated family of possible distributions.
@@
$\mathbb{P}[\sup_{u\,:\,\min(|u-x_{n}|,\,|u+x_{n}|)>c_{1}\epsilon_{N}}\prod_{y\in G_{n,N}^{(k)}}\frac{q_{u}(\delta^{*}_{n,y})}{q_{x_{n}}(\delta^{*}_{n,y})}\geq e^{-c_{2}\ell_{N}\epsilon_{N}^{2}}]\leq 4e^{-c_{3}\ell_{N}\epsilon_{N}^{2}},$ (20)
where the outer probability is taken with respect to the distribution of the data $\{\delta_{ny}\}_{y\in G_{n,N}^{(k)}}$ given $x_{n}$. On the other hand, for all $u$ satisfying $|u-x_{n}|<\frac{1}{\ell_{N}^{3}}$ and all $N$ sufficiently large, we have
$\prod_{y\in G_{n,N}^{(k)}}\frac{q_{x_{n}}(\delta^{*}_{n,y})}{q_{u}(\delta^{*}_{n,y})}\leq 2.$ (21)
@@ A.2 Consistent Estimates of Signs
Let $\hat{\tilde{x}}_{n}$ be the posterior median of the distribution of $|x_{n}|$ given $\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}}$, and similarly for $\hat{\tilde{x}}_{n^{\prime}}$. For $y\in J$, define the Bernoulli random variables $Z_{y}=\textbf{1}_{A_{y}}$, where $A_{y}$ is the event:
$A_{y}=\{\max(\delta_{ny},\delta_{n^{\prime}y})>|\hat{\tilde{x}}_{n}-\hat{\tilde{x}}_{n^{\prime}}|\}.$ (22)
Note that, given $\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}}\cup\{\delta_{n^{\prime}y}\}_{y\in G_{n^{\prime},N}^{(k(n^{\prime}))}},$ the random variables $\{Z_{y}\}_{y\in J}$ are i.i.d. Bernoulli with some parameter $r_{n,n^{\prime}}$. By the same argument as in Lemma 1, we have the posterior concentration bound:
@@
We observe that this will allow us to learn whether $\tilde{x}_{n},\tilde{x}_{n^{\prime}}$ have the same signs (as long as both are far from 0). More precisely, for $y\in J$, define $Y_{y}=\textbf{1}_{B_{y}}$, where
$B_{y}=\{\max(\delta_{ny},\delta_{n^{\prime}y})>|\hat{\tilde{x}}_{n}+\hat{\tilde{x}}_{n^{\prime}}|\}.$ (24)
Then $Y_{y}$ are i.i.d. Bernoulli with some parameter $q_{n,n^{\prime}}$. The following is a direct calculation with Gaussians:  [^1] If $\sigma=0$, we’d just look at the probability that the latent position is in the interval $(-\min(|\hat{\tilde{x}}_{n}|,|\hat{\tilde{x}}_{n^{\prime}}|),\min(|\hat{\tilde{x}}_{n}|,|\hat{\tilde{x}}_{n^{\prime}}|))$, for which this is obvious. Since $\sigma>0$, a complete calculation needs to add in a few additional cases. These doesn’t substantially change the results from the trivial case.

### v3

@@ 3.2 Analysis of global influenza
When incorporated into a larger Bayesian hierarchical model, sBMDS provides dimension reduction that propagates and accounts for total model uncertainty. To demonstrate the robustness and applicability of our sparse frameworks, we integrate them within a Bayesian hierarchical model for analyzing the global spread of influenza. Every year seasonal influenza affects millions of adults, resulting in about 140,000 to 710,000 influenza‐related hospitalizations in the United States alone (Rolfes et al.,, 2018). The virus’s ability to constantly evolve makes understanding its viral patterns so important for managing prevalence. The use of easily accessible mobility data can improve the readiness in which we learn about viral epidemics. Holbrook et al., (2021) apply the BMDS framework to a phylogeographic analysis of the spread of influenza subtypes through transportation networks. They analyze 1,370, 1,389, 1,393 and 1,240 samples of type H1N1, H3N1, Victoria (VIC) and Yamagata (YAM), spanning 12.9, 14.2, 15.4 and 17.75 years, respectively. To scale BMDS to data of this size, they implement core model likelihood and log-likelihood gradient calculations on large graphics processing units and multi-core central processing units. Unfortunately, such an approach requires time-intensive coding and access to expensive computational hardware. We employ a similar Bayesian hierarchical model, applying the same highly structured stochastic process priors but use sBMDS to transform to a latent network space. We are interested in whether under sBMDS we can accurately and efficiently infer the subtype-specific rates of dispersal across the latent airspace for the four influenza strains.
Our data consists of pairwise “effective distances” (Brockmann and Helbing,, 2013) between countries, which inversely measures the probability of traveling between airports. More trafficked airports have a shorter “distance” and thus a higher chance of disease transmission. Effective distances are better at predicting disease arrival times and spread compared to geographical distances because they incorporate the underlying mobility network (Brockmann and Helbing,, 2013). For each influenza subtype, we apply sBMDS to their air traffic data with the following priors for the unknown parameters, $\mathbf{X}_{v},\sigma_{v}^{2}$, and hyperparameter $\mathbf{\Sigma}_{v}$. For strain $v$, the prior on the viral latent locations $\mathbf{X}_{v}$ follows a multivariate Brownian diffusion process along the tree
$\mathbf{X}_{v}\sim MN(\bm{\mu}_{v},\mathbf{V}_{{\cal G}_{v}},\mathbf{\Sigma}_{v}),$ (16)
in which $\bm{\mu}_{v}$ is the $N\times D$ mean matrix, $\mathbf{V}_{{\cal G}_{v}}$ is the $N\times N$ row covariance matrix calculated from a fixed tree ${\cal G}_{v}$, and $\mathbf{\Sigma}_{v}$ is the $D\times D$ column covariance matrix, independently for $v=1,...,4$. For viral diffusion, $\mathbf{\Sigma}_{v}$ describes how the virus’s location in geographic space covary over lineages. In addition, we assume a priori
@@
$d_{0}$ is the degree of freedom set as the dimension of the latent space and $\mathbf{T}_{0}$ is the rate matrix fixed as $\mathbf{I}_{D}$ in our model. The trace of $\mathbf{\Sigma}_{v}$ provides the instantaneous rate of diffusion and is of chief scientific interest. One can think of spatial variance as how much the virus diffuses in a geographic dimension, so by summing up the variance in each dimension, we can understand the total spread of a virus across space in a given moment. We want to accurately infer the trace of $\mathbf{\Sigma}_{v}$ with our phylogenetic sBMDS model trained on a latent airspace. We implement the adaptive HMC algorithm to recover the viral latent locations along with adaptive MH updates on the BMDS precision parameter, $1/\sigma^{2}_{v}$, and Gibbs updates on $\mathbf{\Sigma}^{-1}_{v}$. We let the latent dimension be six as Holbrook et al., (2021) recommended from 5-fold cross-validation. We find 20 leapfrog steps to be adequate as we vary the number of bands/landmarks to 50, 100 and 200.
@@ 3.2.1 Accuracy
For each subtype and model, we run 120,000 iterations, burning the first 20,000 and saving every 100th iteration. Figure 9 plots the posterior distributions of the strain-specific diffusion rates inferred from the full (left) and banded sparse (right) model. We successfully capture the relative distributions for the B-sBMDS using 50 bands, but note that the posterior modes are slightly off. When we increase the number of bands to 200 (Figure B.3), the distributions appear identical. For Figure 10, we apply sBMDS with 50 bands on the H1N1 air traffic data and use procrustes to align the inferred latent locations for each country across all iterations. Since our data has multiple taxon IDs per country, we take the median of the procrustes aligned means of the inferred latent locations and plot the first two dimensions. Figure 10 demonstrates that we obtain a reasonable map; countries in the same continent group together, and within continents, countries with more air traffic are more centrally located. Using the textmineR package (Jones,, 2021) in R, we compute the Hellinger distance between the strain-specific posterior distributions of the squared effective distance per year from the full and sparse methods (Table 3). As expected, the Hellinger distance decreases with more bands.
@@
Figure 10: The first two dimensions of the inferred latent locations for each country using a 6-dimensional sparse Bayesian phylogenetic multidimensional scaling with 50 bands. The plotted inferred latent locations are the median of the Procrustes aligned means across all iterations and taxon.
@@ 3.2.2 Computational performance
B Hellinger distance Average efficiency speedup (min, max)
@@
L Hellinger distance Average efficiency speedup (min, max)
@@
3.3 Cluster analysis of ArXiv articles
We explore the utility of sBMDS for propagating uncertainty in downstream tasks by applying a cluster analysis to a collection of ArXiv paper abstracts. Using the arxivscraper package (Sadjadi,, 2017) in Python, we scrape articles posted on ArXiv from December 2017 to March 2024 across four subject areas: mathematical logic (math.LO), applied physics (physics.app-ph), machine learning (stat.ML) and economics (q-fin.ec). Our final dataset includes 9,308 articles with 1,411 related to math, 1,838 to physics, 5,361 to statistics and 698 to economics. We extract each paper’s abstract and embed it into a 768-dimensional numerical vector using SentenceTransformers (Reimers and Gurevych,, 2019) under the all-mpnet-base-v2 model (Song et al.,, 2020) in Python. This large language model produces sentence-level embeddings that capture semantic similarity; therefore, abstracts with similar content yield similar embeddings. To form an observed dissimilarity matrix, we compute the pairwise cosine dissimilarities between embeddings, e.g, $1-\cos{(\mathbf{x}_{n},\mathbf{x}_{n^{\prime}})}=1-\frac{\mathbf{x}_{n}\cdot\mathbf{x}_{n^{\prime}}}{||\mathbf{x}_{n}||||\mathbf{x}_{n^{\prime}}||}$ for $\mathbf{x}_{n},\mathbf{x}_{n^{\prime}}\in\mathbb{R}^{D}$. Cosine dissimilarity is appropriate for textual data, as it emphasizes directional similarity over magnitude.
We apply B-sBMDS to this 9,308 by 9,308 observed distance matrix using 50, 100, 500, 1,000 and 2,000 bands and obtain the posterior distribution of the latent locations over a 2-dimensional space. We use an adaptive HMC algorithm with 20 leapfrog steps to recover the latent locations. We keep every 100th iteration and run enough iterations such that the ESS approximates the number of thinned samples, indicating “near-independence” among samples. To incorporate the uncertainty encoded in the sBMDS posterior samples, we implement a “bagged estimator”-style algorithm. We randomly draw $S$ iterations from the joint posterior distribution of the latent locations and implement hierarchical density-based spatial clustering of applications with noise (HDBSCAN) for each iteration using the dbscan package (Hahsler et al.,, 2019) in R. DBSCAN (Ester et al.,, 1996) is a non-parametric clustering algorithm that groups points into dense regions based on a user-defined radius parameter $\epsilon$ and a minimum number of neighbors. Points within a dense region are assigned to the same cluster, while points in low-density regions are labeled as noise. HDBSCAN (Campello et al.,, 2015) performs DBSCAN for various $\epsilon$ values and integrates the results to give the most stable output.
Figure 11: Convex cluster hulls for five posterior samples of the inferred latent locations using hierarchical density-based spatial clustering (HDBSCAN). The inferred locations are estimated from a 2-dimensional B-sBMDS framework using 500 bands and an adaptive Hamiltonian Monte Carlo (HMC) algorithm. The convex hulls are plotted on the posterior mean of the inferred locations across 1,000 iterations, and points outside a hull are considered noise. The ArXiv subject of a cluster is assumed by the most frequent ArXiv category among its members.
Figure 11 visualizes the results of HDBSCAN applied to five posterior samples of the latent space plotted on the posterior mean of the latent locations. We set the minimum number of neighbors to 40. Table 4 presents clustering accuracy across band settings. Each cluster is assigned a subject label based on the most frequent ArXiv category among its members. Classification error is the proportion of articles whose ArXiv subject label does not match the subject label of their assigned cluster. We calculate the mean classification error across clusters and note that 500 bands provide a good balance between accuracy and computational cost.
Band Mode number of clusters Mean number of noise points Mean classification error (95% CI)
50 2 3975 0.226 (0.111, 0.315)
100 3 2954 0.094 (0.038, 0.171)
500 4 1380 0.051 (0.041, 0.061)
1000 4 1248 0.038 (0.031, 0.045)
2000 4 1009 0.034 (0.031, 0.037)
Table 4: Accuracy as the number of bands increases using hierarchical density-based spatial clustering (HDBSCAN) on 200 randomly sampled inferred latent locations. The inferred locations are estimated from a 2-dimensional B-sBMDS framework and an adaptive Hamiltonian Monte Carlo (HMC) algorithm. Note that the number of clusters should be 4. HDBSCAN is a non-parametric clustering algorithm and therefore the number of clusters is predicted. Classification error is the proportion of articles whose ArXiv subject label does not match the majority subject label of their assigned cluster, and CI is the credible interval.
We obtain clustering labels for each article indirectly induced by sBMDS’s uncertainty. Across the $S$ iterations, we compute a co-clustering matrix that estimates the posterior pairwise probability that two articles belong in the same cluster. This matrix can be used for subsequent analyses such as consensus clustering. The combination of sBMDS with HDBSCAN provides an example of how to propagate uncertainty from a sparse Bayesian embedding model to post-hoc inference.
@@ 4 Discussion
We present two methods for subsetting the observed dissimilarity data: banded sparse BMDS (B-sBMDS) and landmark sparse BMDS (L-sBMDS). We show that both sparse methods obtain accurate results at low band/landmark counts even with noisy data. Moreover, combining HMC with sBMDS proves effective in inferring thousands of latent locations. We successfully integrate the sBMDS variants within a Bayesian hierarchical model and demonstrate that one may propagate uncertainty represented by the sBMDS posterior to downstream modeling tasks. In well-specified cases, we recommend using L-sBMDS, as it achieves higher accuracy with fewer landmarks, leading to greater speedups compared to B-sBMDS. However, in most cases, B-sBMDS is the better choice due to its robustness to model misspecification. For example, in our influenza application, in which the Euclidean assumption is violated, B-sBMDS yields a slightly smaller Hellinger distance than L-sBMDS.
Possible extensions to our work include the use of different noise distributions on the observed dissimilarities. For example, Bakker and Poole, (2013) employ Bayesian metric MDS, assuming the observed dissimilarities come from log-normal distributions. As these distributions still have $\mathcal{O}(N^{2})$ time complexity, the sBMDS could be valuable in improving the computational performance for a wider range of dissimilarity data.
Additionally, many potential theoretical developments remain. For example, it appears that one needs approximately $D\sqrt{N}$ bands where $N$ is the number of objects and $D$ is the embedding space. However, we have no formal proof, only experimental results (Sections 3.2 and 3.3, Figure B.4). Determining the number of desired bands/landmarks is difficult due to its data-dependence. We explain in Section A.4 how one could extend Theorem 1’s proof of posterior consistency to higher dimensions. The biggest limitations are extending Lemma 3 and obtaining estimates with good dependence on dimension $D$. One could also explore treating the coupling matrix $J_{n,N}$ as a random variable that depends on the observed data (and perhaps changes over the run-time of an algorithm). An appealing feature of Raftery et al., (2012) is that they claim reasonable uncertainty quantification along a truly linear run-time. It seems difficult to formalize such a result with posterior consistency for our sBMDS models as the number of bands (landmarks) grows with the number of objects. We are left with many tantalizing questions: “by including a data-informed approach to model sparsity, can we achieve a linear run-time and still demonstrate posterior consistency?”, “how should we be measuring consistency?”, and “do the datasets Raftery et al., (2012) study have any special features that change the rate of convergence for a sBMDS-like model?”
@@ A.1 Consistent Estimates of Absolute Values
We note that $|x_{n}|$ (but not $x_{n}$ itself) is effectively identifiable given the data $\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}}$, and we have the posterior concentration bound:
@@
$\mathbb{P}[p(\{u\,:\,\min(|u-x_{n}|,|u+x_{n}|)\leq c_{1}\epsilon_{N}\}|\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}})\geq 1-e^{-c_{2}\ell_{N}\epsilon_{N}^{2}}]\geq 1-e^{-c_{3}\ell_{N}\epsilon_{N}^{2}}.$ (19)
@@
Given $x_{n}$, the data $\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}}$ are i.i.d. with distributions being a finite mixture of truncated Gaussians. Denote the density of this distribution by $q_{x_{n}}$, and let =$\mathcal{F}=\{q_{u}\}_{u\in I}$ be the associated family of possible distributions.
@@
$\mathbb{P}[\sup_{u\,:\,\min(|u-x_{n}|,\,|u+x_{n}|)>c_{1}\epsilon_{N}}\prod_{y\in G_{n,N}^{(k(n))}}\frac{q_{u}(\delta^{*}_{n,y})}{q_{x_{n}}(\delta^{*}_{n,y})}\geq e^{-c_{2}\ell_{N}\epsilon_{N}^{2}}]\leq 4e^{-c_{3}\ell_{N}\epsilon_{N}^{2}},$ (20)
where the outer probability is taken with respect to the distribution of the data $\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}}$ given $x_{n}$. On the other hand, for all $u$ satisfying $|u-x_{n}|<\frac{1}{\ell_{N}^{3}}$ and all $N$ sufficiently large, we have
$\prod_{y\in G_{n,N}^{(k(n))}}\frac{q_{x_{n}}(\delta^{*}_{n,y})}{q_{u}(\delta^{*}_{n,y})}\leq 2.$ (21)
@@ A.2 Consistent Estimates of Signs
Let $\hat{\tilde{x}}_{n}$ be the posterior median of the distribution of $|x_{n}|$ given $\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}}$, and similarly for $\hat{\tilde{x}}_{n^{\prime}}$. For $j\in J$, define the Bernoulli random variables $Z_{j}=\textbf{1}_{A_{j}}$, where $A_{j}$ is the event:
$A_{j}=\{\max(\delta_{nj},\delta_{n^{\prime}j})>|\hat{\tilde{x}}_{n}-\hat{\tilde{x}}_{n^{\prime}}|\}.$ (22)
Note that $\hat{\tilde{x}}_{n},\hat{\tilde{x}}_{n^{\prime}}$ are $\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}}\cup\{\delta_{n^{\prime}y}\}_{y\in G_{n^{\prime},N}^{(k(n^{\prime}))}}$-measurable, and $(\delta_{nj},\delta_{n^{\prime}j})$ are independent of $\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}}\cup\{\delta_{n^{\prime}y}\}_{y\in G_{n^{\prime},N}^{(k(n^{\prime}))}}$ for each $j\in J$, and finally the collection $\{(\delta_{nj},\delta_{n^{\prime}j})\}_{j\in J}$ are independent. Thus, conditional on $\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}}\cup\{\delta_{n^{\prime}y}\}_{y\in G_{n^{\prime},N}^{(k(n^{\prime}))}},$ the random variables $\{Z_{y}\}_{y\in J}$ are i.i.d. Denote by $r_{n,n^{\prime}}$ their common parameter. By the same argument as in Lemma 1, we have the posterior concentration bound:
@@
We observe that this will allow us to learn whether $\tilde{x}_{n},\tilde{x}_{n^{\prime}}$ have the same signs (as long as both are far from 0). More precisely, for $j\in J$, define $Y_{j}=\textbf{1}_{B_{j}}$, where
$B_{j}=\{\max(\delta_{nj},\delta_{n^{\prime}j})>|\hat{\tilde{x}}_{n}+\hat{\tilde{x}}_{n^{\prime}}|\}.$ (24)
By the same argument as the one immediately following Equation (22), the $Y_{j}$ are i.i.d. Bernoulli. Denote by $q_{n,n^{\prime}}$ their common parameter. The following is a direct calculation with Gaussians [^1] If $\sigma=0$, we’d just look at the probability that the latent position is in the interval $(-\min(|\hat{\tilde{x}}_{n}|,|\hat{\tilde{x}}_{n^{\prime}}|),\min(|\hat{\tilde{x}}_{n}|,|\hat{\tilde{x}}_{n^{\prime}}|))$, for which this is obvious. Since $\sigma>0$, a complete calculation needs to add in a few additional cases. These doesn’t substantially change the results from the trivial case.: